corporate investments The financial board of a corporation invested $25000 overseas, part at 4% and part at 7% annual interest.Find the amount invested at each rate if the first year combined income from the two investment was $1300
Found 2 solutions by Theo, mananth:
Answer by Theo(13342)

x = Amount of money invested at 4%
y = Amount of money invested at 7%

Combined Interest = 1300

Combined Interest Equation equals:

.04*x + .07*y = 1300

Total invested = 25,000

Total Investment Equation equals:

x + y = 25000

You need to solve these two equations simultaneously to get your answer.

The equations are:

x + y = 25000
.04x + .07y = 1300

use the first equation to solve for y in terms of x.

you get:

y = 25000 - x

substitute for y in the second equation to get:

.04x + .07(25000-x) = 1300

simplify to get:

.04x + .07*25000 - .07*x = 1300

simplify further to get:

.04x + 1750 - .07x = 1300

combine like terms to get:

-.03x + 1750 = 1300

subtract 1750 from both sides of equation to get:

-.03x = 1300 - 1750 = -450

divide both sides of equation by -.03 to get:

x = -450/-.03 = 150000

since x + y = 2500, then y = 10000

your answer should be:

Amount invested at 4% = 15000.

Amount invested at 7% = 10000.

Interest on 15000 at 4% = .04 * 15000 = 600.

Interest on 10000 at 7% = .07 * 10000 = 700.

Combined interest = 600 + 700 = 1300, confirming the answer is good.






Answer by mananth(16949)

corporate investments The financial board of a corporation invested $25000 overseas, part at 4% and part at 7% annual interest.Find the amount invested at each rate if the first year combined income from the two investment was $1300

Amount invested be x in 7% deposit
Amount invested in 4% deposit will be 25000-x
the iinterest earned on the first one above will be 0.07x
The interest earned on the second one will be 0.04(25000-x)
The sum of both the interests is 1300

0.07x + 0.04(25000-x)= 1300
0.07x + 1000 - 0.04x = 1300
0.07x-0.04x= 1300 - 1000
0.03x = 300
x= 10000
26000-x=25000-10000=15000

$15000 was invested at 4% and $10000 at 7%